OpenStudy (anonymous):

2. Find the coordinates of a point symmetric to A:(3,2) with respect to the line ? can you help me get started?

OpenStudy (anonymous):

It should read Find the coordinates of a point symmetric to A:(3,2) with respect to the line y=3x-2

OpenStudy (anonymous):

it means u need the image of the point A:(3,2) when y=3x-2 acts as mirror

OpenStudy (anonymous):

first find a line perpendicular to y=3x -2 and passing throgh (3,2) which is y -2 =-1/3 (x-3)

OpenStudy (anonymous):

or x+3y=9 now let us find the pt of intersection for x+3y=9 and the given line 3x-y=2 now the point of intersection will give u the midpoint of the required point and (3,2) and hence the reqd pt can be calculated